### v1

# When Reasoning Meets Compression: Benchmarking Compressed Large Reasoning Models on Complex Reasoning Tasks

Abstract
Recent open-source large reasoning models (LRMs) exhibit strong performance on complex reasoning tasks, but their large parameter count makes them prohibitively expensive for individuals. The compression of large language models (LLMs) offers an effective solution to reduce cost of computational resources. However, systematic studies on the performance of compressed LLMs in complex reasoning tasks, especially for LRMs, are lacking. Most works on quantization and pruning focus on preserving language modeling performance, while existing distillation works do not comprehensively benchmark student models based on reasoning difficulty or compression impact on knowledge and reasoning. In this paper, we benchmark compressed DeepSeek-R1 models on four different reasoning datasets (AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue), ranging from mathematical to multihop reasoning, using quantization, distillation, and pruning methods. We benchmark 2.51-, 1.73-, and 1.58-bit R1 models that adopt dynamic quantization. We also benchmark distilled R1 models that are based on LLaMA or Qwen and run SparseGPT on them to obtain various sparsity levels. Studying the performance and behavior of compressed LRMs, we report their performance scores and test-time compute (number of tokens spent on each question). Notably, using MuSiQue, we find that parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability, which can inform the choice of compression techniques. Through our empirical analysis of test-time compute, we find that shorter model outputs generally achieve better performance than longer ones across several benchmarks for both R1 and its compressed variants, highlighting the need for more concise reasoning chains.
1 Introduction
Large reasoning models (LRMs) excel at complex reasoning tasks. DeepSeek-R1 (Guo et al., 2025) is the first open-source LRM to successfully replicate OpenAI’s O1-level performance. Trained based on DeepSeek-V3 (DeepSeek-AI et al., 2025), R1 adopts large-scale reinforcement learning without supervised fine-tuning (SFT). Its open-source nature facilitates research on reasoning models (Srivastava et al., 2025; Mondillo et al., 2025). However, due to its large size (671B total parameters), deploying it locally can be costly and even infeasible for individuals, which hinders AI democratization.
Compression of large language models (LLMs) includes quantization, distillation, and pruning. Its goal is to reduce computational resources (e.g., GPU memory and disk space) and provide inference speedup. As the weights of vanilla LLMs are typically in 16-bit floating values (e.g.., FP16 or BF16), quantization means to convert these high-precision values to lower-precision ones. Representative techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization (Lin et al., 2024), and post-training quantization (Frantar et al., 2022). Distillation aims to transfer knowledge from a larger teacher model to a smaller student model under black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings. As for pruning, it aims to sparsify LLMs by introducing 0s to replace nonzero weights. Representative techniques include unstructured pruning (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
Although existing quantization and pruning methods claim to preserve the performance of non-reasoning LLMs after compression such as LLaMA (Touvron et al., 2023) and OPT (Zhang et al., 2022)), exploring these methods on LRMs offers a timely research of democratizing the latest LRMs. Moreover, since these methods mainly adopt perplexity and simple end-task scores for evaluation, the assessment of compressed LRMs will be fundamentally different. Without focusing on basic linguistic capabilities (e.g., simple language understanding and generation), evaluation of compressed reasoning models needs to align with their uncompressed counterparts. Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider compression impact on knowledge and reasoning. A comprehensive evaluation that includes all popular compression methods is valuable.
Therefore, due to the lack of compression research on LRMs, we benchmark the performance and behavior of compressed DeepSeek-R1 on various reasoning tasks in this paper. Our benchmarking framework is shown in Figure 1. We test dynamic quantization (Daniel Han & team, 2023), distillation with SFT (Guo et al., 2025), and SpraseGPT (Frantar & Alistarh, 2023) on R1 (or distilled R1) due to their popularity and effectiveness. Specifically, our chosen dynamic quantization reduces the MoE (Mixture of Experts) layers of R1 model to 2.51-bit, 1.73-bit, and 1.58-bit precision. Testing distillation with SFT, we evaluate the distilled R1 models that adopt LLaMA-70B (Meta AI, ) and Qwen-32B (Qwen et al., 2025) as students. Using SparseGPT, we prune the distilled R1 models to different sparsity levels. We select four complex reasoning benchmarks: mathematical, logical, temporal, and multihop reasoning.
Figure 1: An overview of selected compression methods and tasks. We evaluate R1 and its compressed variants based on their performance scores and test-time compute.
In addition to assessing model performance on selected benchmarks, we study the behavior of compressed R1 models by measuring their test-time compute (number of tokens spent on each question). Specifically, we compare their performance on the shortest and longest outputs. Our major findings are:
  • Parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability. Therefore, both distillation and pruning are discouraged when the end task relies on models’ parametric knowledge. Otherwise, larger student models or low sparsity levels are recommended.
  • In almost all cases, both R1 and its compressed variants achieve higher scores when they spend less compute during inference time. Shorter outputs tend to be more accurate than longer ones across our benchmarks. This highlights the need to reduce verbosity to improve reasoning performance.
2 Related Work
2.1 Quantization
Quantization reduces the number of bits used to represent LLM weights, thereby lowering their precision (Srivastava et al., 2025). Recent survey (Zhu et al., 2024) categorizes quantization methodologies into quantization-aware training (QAT) and post-training quantization (PTQ). QAT requires retraining of model weights to recover performance loss during quantization while PTQ does not require retraining. Recent QAT includes LLM-QAT (Liu et al., 2024a) that adopts distillation to train a quantized LLM, BitDistiller (Du et al., 2024) that develops a self-distillation approach for the full-precision model to act as the teacher of its low-bit counterpart, BitNet (Wang et al., 2023) that proposes a 1-bit Transformer architecture for training LLMs from scratch, and OneBit (Xu et al., 2024) that quantizes LLM weight matrices to 1-bit from a knowledge transfer perspective.
PTQ is more popular in terms of the number of recent publications, because there is no retraining involved. For example, GPTQ (Frantar et al., 2022) is a one-shot weight quantization method that uses approximate second-order information, while AWQ (Lin et al., 2024) leverages activation distribution for finding the salient weight channels to skip. Other PTQ methods include weight-activation quantization (Shao et al., 2024; Yao et al., 2022; Liu et al., 2023) and KV cache quantization (Hooper et al., 2024; Liu et al., 2024b).
2.2 Distillation
Distillation involves two settings: black-box and white-box settings. For black-box setting, teacher model is typically a closed-source LLM and only the outputs of teacher are available for student model. For white-box setting, both weights and output distribution of the teacher model are available. Existing black-box distillation (Huang et al., 2024; Li et al., 2024b; Ho et al., 2023; Huang et al., 2022; Li et al., 2024a) prompts the teacher model to generate a training dataset for the student to learn. Specifically, researchers have started to distill OpenAI’s O1 model (Huang et al., 2024), which marks the beginning of LRMs compression. White-box distillation allows the student model to learn from teacher’s knowledge representation. Works has been done to align the output distribution (Agarwal et al., 2024; Gu et al., 2024) or the hidden representation (Liang et al., 2023) between teacher and student models.
2.3 Pruning
There are unstructured and structured pruning. For unstructured pruning, individual weights are targeted, which leads to irregular sparsity structure. In contrast, structured pruning involves removing entire network components such as channels or layers (Zhang et al., 2024a). Unstructured pruning usually has better compression performance than structured pruning, while it is easier to achieve inference speedup via structured methods (Zhu et al., 2024). Recent unstructured pruning includes one-shot pruning (Frantar & Alistarh, 2023; Sun et al., 2023), global pruning that makes pruning decisions based on all layers (Bai et al., 2024), and domain-specific pruning (Zhang et al., 2024a). Structured pruning includes gradient-based  (Xia et al., 2024; Ma et al., 2023) and non-gradient-based (Ashkboos et al., 2024) methods.
2.4 LRMs
Trained with reinforcement learning, LRMs extends LLMs with advanced reasoning mechanisms (Besta et al., 2025). Popular closed-source LRMs are OpenAI’s o1-mini, o1 (OpenAI et al., 2024), and o3-mini. Open-source LRMs include DeepSeek-R1 and QwQ-32B-Preview (Team, 2024). Since quantization, white-box distillation, and pruning methods require access to model weights, they are not suitable for closed-source LRMs. Only black-box distillation will work on closed-source models.
3 Background
Scope We benchmark quantization, distillation, and pruning methodologies, because we consider them as major LLMs compression paradigms in this paper. Due to lower popularity, other methodologies that can potentially facilitate compression such as low-rank factorization (Sharma et al., 2024; Srebro & Jaakkola, 2003) are not elaborated. When selecting models for compression, we focus exclusively on LRMs.
Recent efforts on LRMs compression As reviewed in Section 2, compression on LRMs (not LLMs) is relatively underexplored. Recently, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights. By combining four ideas (Yu et al., 2024; Dettmers et al., 2023; Ma et al., 2024; Gerganov, 2023), Unsloth quantizes DeepSeek-R1 into four variants: 2.51-, 2.22-, 1.73-, and 1.58-bit models. Here the number of bits (e.g., 2.51-bit) represents the precision of MoE weights, which make up 88% of the total R1 weights. DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation.
Current bottlenecks Very few quantization or pruning works have demonstrated effectiveness on LRMs, as LRMs have only recently emerged. Moreover, evaluation of quantized or pruned LRMs will be different from quantized or pruned LLMs. Current works evaluate quantization and pruning performance primarily using perplexity and simple end tasks, such as the EleutherAI evaluation harness (Gao et al., 2024) and commonsense reasoning. However, compressed LRMs should be assessed on more complex reasoning tasks with varying difficulty levels. As for distillation, although recent distillation works tend to test on more challenging reasoning tasks (compared to other compression literature) such as GSM8K (Cobbe et al., 2021), it is unclear how the compression of LRMs affects models’ parametric knowledge and reasoning capability. Some of them do not comprehensively select diverse reasoning benchmarks. Our benchmarking framework aims to address these bottlenecks.
#Questions #Pass Answer Type Metric Knowledge required?
AIME 2024 30 1 Integer Accuracy False
FOLIO 203 1 True/False/Uncertain Accuracy False
Temporal 250 1 (A)/(B)/(C)/(D) Accuracy False
MuSiQue 100 1 A few words (EM, F1) True
Table 1: Statistics of selected benchmarks: AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue. We use exact match (EM) and F1 to measure model performance on MuSiQue.
4 Experiment Setup
4.1 Reasoning Benchmarks
To address the evaluation bottleneck discussed in Section 3, we select four different reasoning benchmarks with varying levels of difficulty: AIME 2024 (Mathematical Association of America, ) for mathematical reasoning, FOLIO (Han et al., 2024) for logical reasoning, Temporal Sequences of BIG-Bench Hard (Suzgun et al., 2022) for temporal reasoning, and MuSiQue (Trivedi et al., 2022) for multihop reasoning. Table 1 summarizes their statistics. For each model, we do a single pass on every benchmark.
Unlike other benchmarks where we select all questions, we randomly sample 100 out of 1000 from MuSiQue, as it is much larger. This pool of 1000 questions is the same as SiReRAG (Zhang et al., 2025) and HippoRAG (Jimenez Gutierrez et al., 2024). Since MuSiQue requires knowledge memorization besides multihop reasoning, it is popular under retrieval-augmented generation (RAG) setup. Instead of RAG, we follow a closed-book setting (directly prompting LRMs to get final answers) to test models’ parametric knowledge. In this case, unlike other benchmarks that test only reasoning capability, we use MuSiQue to evaluate both reasoning and knowledge memorization. Additional details of selected benchmarks are specified in Appendix A.
4.2 Selected Compression Methods
We decide to compress DeepSeek-R1 due to it large parameter count. For quantization, we select 2.51-, 1.73-, and 1.58-bit models [^1] https://huggingface.co/unsloth/DeepSeek-R1-GGUF by Unsloth. These dynamically quantized models are among the most widely used quantization methods on R1. As for distillation, we select the two largest distilled models that accompany R1: DeepSeek-R1-Distill-Llama-70B [^2] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-32B [^3] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-32B. Since we do not find a pruning method designed for models with around 700B parameters and MoE layers (e.g., SparseGPT can only prune models with around 176B parameters), we use SparseGPT to prune our selected distilled models. In order to precisely analyze the collapse point of this pruning method, we run SparseGPT on each distilled model seven times to get 10%, 30%, 40%, 50%, 60%, 70%, and 80% sparsity levels.
4.3 Evaluation Metrics
Accuracy metric is used for AIME 2024, FOLIO, and Temporal Sequences. Since MuSiQue involves question answering and its answers are in a few words, we adopt exact match (EM) and F1 to measure question answering performance of different models. Without worrying inference speedup of selected compression methods, we focus on benchmarking their performance. The reason is that these methods run on different inference platforms. As a result, different optimization strategies might be adopted when we do inference, and it is hard to control the consistency of inference optimization across various platforms.
Our analysis of test-time compute involves computing the number of tokens spent on each question. We first preprocess each output by extracting reasoning tokens and sentences of final model prediction. Instead of relying on a specific tokenizer, we then treat words within each output as tokens for simplicity. We study whether there is any relationship between length and reasoning performance.
4.4 Implementation Details
Required by our quantized models, we run their inference on llama.cpp [^4] https://github.com/ggml-org/llama.cpp. While our pruned and distilled models can be deployed on various inference platforms, we use vLLM (Kwon et al., 2023) for its fast inference. In order to comprehensively analyze performance change after compression, we also evaluate R1 on our reasoning benchmarks by using DeepSeek API. Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models: maximum generation length is set to $3276832768$, temperature is set to $0.60.6$, and top-p value is set to $0.950.95$.
5 Results and Analysis
Our results and analysis aim to answer the following research questions:
  • RQ 1: How does each compression methodology compare against each other (5.1)?
  • RQ 2: Do our selected compression methodologies collapse (5.2)?
  • RQ 3: What is the recommended methodology to compress LRMs (5.1, 5.3, and 5.5)?
  • RQ 4: Is it a good sign when models spend extra amounts of test-time compute on certain questions (5.4)?
Models Accuracy
Model #Param Compression AIME 2024 FOLIO Temporal Avg MuSiQue (EM, F1)
DeepSeek-R1 671B - 73.3 76.4 99.6 83.1 (17.0, 27.51)
DeepSeek-R1 671B 2.51-bit 76.7 77.8 100.0 84.8 (17.0, 24.43)
DeepSeek-R1 671B 1.73-bit 66.7 78.3 99.6 81.5 (15.0, 22.11)
DeepSeek-R1 671B 1.58-bit 66.7 75.4 94.0 78.7 (14.0, 22.34)
R1-Distill-Llama 70B Distillation 63.3 78.8 100.0 80.7 (13.0, 21.80)
R1-Distill-Llama 70B Distillation & 50% sparse 26.7 70.9 97.2 64.9 (6.0, 12.75)
R1-Distill-Qwen 32B Distillation 66.7 82.3 100.0 83.0 (1.0, 9.38)
R1-Distill-Qwen 32B Distillation & 50% sparse 30.0 75.4 96.0 67.1 (3.0, 9.29)
Table 2: Overall results of R1 and it compressed variants. The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal. EM and F1 of MuSiQue are shown in tuples. We segment this table based on model family (e.g., R1-Distill-Llama).
5.1 Overall Performance
The overall results of R1 and its compressed variants are presented in Table 2 and 3. We only show performance of pruned R1-Distill-Llama and pruned R1-Distill-Qwen under 50% sparsity in Table 2, as it is the default sparsity level of SparseGPT and many other works (Zhang et al., 2024a; Sun et al., 2023). Table 3 displays performance of our two distilled models under various sparsity levels.
Comparing Performance of Compression Methodologies In Table 2, 2.51-bit R1 achieves the highest average accuracy scores, surpassing even the original R1. We suspect that certain security mechanisms in DeepSeek may have lowered R1’s performance, as we often receive empty responses when using its API, requiring us to prompt R1 iteratively until valid responses are obtained. Both R1-Distill-Llama and R1-Distill-Qwen have competitive average accuracy scores. According to Table 3, we see that pruning also maintains good average accuracy scores when sparsity levels are low (e.g., 30% sparsity of R1-Distill-Llama). On MuSiQue dataset, the 2.51-bit model has the same EM as R1, with a slight decrease in F1. According to scores of the closed-book setting on GPT-4o (Zhang et al., 2025), all quantized models achieve relatively strong performance on MuSiQue, with higher EM and F1 than distilled and pruned models. Therefore, this 2.51-bit model has the best overall performance than other compression methods.
Comparing Benchmark Difficulties We can make judgment about the difficulties of our accuracy-based benchmarks through Table 2. By comparing the scores of AIME 2024, FOLIO, and Temporal, we see all models struggle more on AIME 2024. This indicates that AIME 2024 is more difficult than the other two benchmarks. MuSiQue is also a difficult benchmark in terms of knowledge requirement, because its scores in Table 2 are much lower than RAG setup (Zhang et al., 2025; Jimenez Gutierrez et al., 2024). This suggests that existing LRMs lack sufficient knowledge for knowledge-intensive tasks, making RAG a more suitable approach. We demonstrate a diverse selection of reasoning benchmarks with varying difficulty levels, which is important for comprehensive evaluation.
Comparing Distilled Models based on LLaMA and Qwen When focusing on accuracy-based benchmarks, we see that R1-Distill-Qwen delivers an average 2.3% improvement over R1-Distill-Llama. This demonstrates that Qwen-32B has stronger reasoning capability than LLaMA-70B, which also aligns with the evaluation results of DeepSeek-R1 report (Guo et al., 2025). However, R1-Distill-Qwen scores significantly lower on MuSiQue, highlighting its inability to memorize detailed knowledge.
5.2 Collapse Point of Each Compression Method
We investigate whether LRMs degrade as they undergo increasing levels of compression. In Table 2, the performance of dynamically quantized models steadily declines as we move from 2.51-bit to 1.58-bit. However, we do not observe a clear collapse point, as performance declines only slightly when transitioning to a lower bit-size tier. As for distillation, comparing R1 with R1-Distill-Llama, we also do not observe a clear collapse point, although the performance drop on MuSiQue is significant. However, we see that R1-Distill-Qwen collapses on MuSiQue. With EM close to zero, it almost loses its ability on MuSiQue during distillation.
Comparing distilled models with their sparsified variants in Table 3, we find that pruned models collapse on all benchmarks at certain sparsity levels. Interestingly, their collapse point correlates to the difficulty of the benchmark. For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half. However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024. The correlation between collapse point and benchmark difficulty can also be seen on sparsified R1-Distill-Qwen. The early occurrence of collapse on AIME 2024 demonstrates that it is one of the most challenging datasets, as existing pruning works typically observe model collapse after 50% sparsity (Zhang et al., 2024a; Bai et al., 2024). Moreover, since R1-Distill-Qwen already collapses on MuSiQue, pruning it becomes meaningless for this dataset, as we observe very low scores for all sparsified R1-Distill-Qwen models on MuSiQue.
Models Accuracy
Model #Param Sparsity AIME 2024 FOLIO Temporal Avg MuSiQue (EM, F1)
R1-Distill-Llama 70B 0% 63.3 78.8 100.0 80.7 (13.0, 21.80)
R1-Distill-Llama 70B 10% 60.0 81.3 99.6 80.3 (12.0, 21.69)
R1-Distill-Llama 70B 30% 63.3 79.3 99.6 80.7 (14.0, 21.40)
R1-Distill-Llama 70B 40% 56.7 73.9 98.8 76.8 (6.0, 13.79)
R1-Distill-Llama 70B 50% 26.7 70.9 97.2 64.9 (6.0, 12.75)
R1-Distill-Llama 70B 60% 0.0 65.0 95.6 53.5 (0.0, 6.42)
R1-Distill-Llama 70B 70% 0.0 49.8 15.6 21.8 (0.0, 2.23)
R1-Distill-Llama 70B 80% 0.0 11.8 12.4 8.1 (0.0, 0.94)
R1-Distill-Qwen 32B 0% 66.7 82.3 100.0 83.0 (1.0, 9.38)
R1-Distill-Qwen 32B 10% 70.0 81.3 100.0 83.8 (5.0, 13.19)
R1-Distill-Qwen 32B 30% 56.7 81.3 100.0 79.3 (1.0, 10.47)
R1-Distill-Qwen 32B 40% 53.3 78.3 100.0 77.2 (2.0, 10.16)
R1-Distill-Qwen 32B 50% 30.0 75.4 96.0 67.1 (3.0, 9.29)
R1-Distill-Qwen 32B 60% 0.0 65.0 87.2 50.7 (0.0, 4.13)
R1-Distill-Qwen 32B 70% 0.0 32.5 19.6 17.4 (0.0, 1.72)
R1-Distill-Qwen 32B 80% 0.0 8.7 2.0 3.6 (0.0, 1.29)
Table 3: Performance of our two distilled models under various sparsity levels. The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal.
Models AIME 2024 FOLIO MuSiQue
Model Compression Short Long Ratio Short Long Ratio Short Long Ratio
DeepSeek-R1 - 88.9 33.3 5.3 83.3 63.3 8.0 (30.0, 42.8) (3.3, 10.0) 6.4
DeepSeek-R1 2.51-bit 100.0 33.3 4.9 85.0 71.7 6.5 (33.3, 41.9) (0.0, 6.9) 7.7
DeepSeek-R1 1.73-bit 88.9 22.2 4.4 86.7 73.3 4.9 (30.0, 41.6) (10.0, 23.5) 6.9
DeepSeek-R1 1.58-bit 77.8 44.4 3.8 80.0 65.0 5.0 (30.0, 41.5) (10.0, 16.6) 5.9
R1-Distill-Llama Distillation 88.9 11.1 5.4 80.0 80.0 4.5 (26.7, 38.0) (6.7, 18.4) 4.0
R1-Distill-Llama Distillation & 10% sparse 100.0 0.0 6.6 85.0 78.3 4.9 (20.0, 29.6) (6.7, 13.7) 7.4
R1-Distill-Llama Distillation & 30% sparse 88.9 11.1 5.3 85.0 76.7 5.0 (26.7, 36.9) (3.3, 12.5) 8.8
R1-Distill-Qwen Distillation 100.0 11.1 7.1 86.7 75.0 6.9 (16.7, 24.1) (16.7, 24.7) 7.1
R1-Distill-Qwen Distillation & 10% sparse 88.9 33.3 4.8 83.3 75.0 5.5 (23.3, 36.7) (3.3, 12.8) 8.1
R1-Distill-Qwen Distillation & 30% sparse 88.9 11.1 6.8 86.7 78.3 7.8 (26.7, 40.7) (3.3, 9.1) 8.8
Table 4: Analysis of test-time compute when selecting the shortest and longest 30% of responses output by each model on each benchmark. “Short” column contains performance scores of the shortest 30% of outputs from a model, while “long” column contains scores of the longest 30% of outputs. We compare the scores between “Short” and “Long” for every model and benchmark, and mark the best scores in bold. “Ratio” column represents the ratio of the average length (in number of tokens) of the longest 30% to that of the shortest 30%.
Models AIME 2024 FOLIO MuSiQue
Model Compression Short Long Ratio Short Long Ratio Short Long Ratio
DeepSeek-R1 - 100.0 16.7 7.1 82.5 57.5 10.9 (30.0, 42.4) (5.0, 12.5) 8.5
DeepSeek-R1 2.51-bit 100.0 33.3 6.6 87.5 62.5 8.6 (40.0, 50.3) (0.0, 4.5) 10.7
DeepSeek-R1 1.73-bit 100.0 33.3 5.9 87.5 67.5 6.3 (30.0, 43.7) (5.0, 15.7) 9.3
DeepSeek-R1 1.58-bit 100.0 33.3 5.6 80.0 72.5 6.2 (35.0, 46.3) (15.0, 20.0) 7.6
R1-Distill-Llama Distillation 83.3 16.7 7.0 82.5 75.0 5.7 (35.0, 45.2) (5.0, 9.2) 5.3
R1-Distill-Llama Distillation & 10% sparse 100.0 0.0 9.2 90.0 75.0 6.3 (20.0, 28.5) (5.0, 11.7) 10.7
R1-Distill-Llama Distillation & 30% sparse 100.0 16.7 7.1 90.0 75.0 6.4 (20.0, 31.0) (0.0, 7.3) 13.0
R1-Distill-Qwen Distillation 100.0 0.0 9.8 87.5 75.0 9.0 (20.0, 28.7) (20.0, 24.2) 10.4
R1-Distill-Qwen Distillation & 10% sparse 83.3 33.3 6.1 85.0 75.0 7.2 (30.0, 36.2) (0.0, 8.4) 12.0
R1-Distill-Qwen Distillation & 30% sparse 100.0 0.0 8.9 87.5 75.0 11.0 (30.0, 36.6) (5.0, 11.2) 13.4
Table 5: Analysis of test-time compute when selecting the shortest and longest 20% of responses output by each model on each benchmark. Refer to Table 4 for the meaning of each column, except that we select shortest and longest 20% instead.
5.3 Compression Impact on Knowledge and Reasoning
As discussed above, although Qwen-32B shows stronger reasoning capability than LLaMA-70B, it has significantly lower EM and F1 scores on MuSiQue. Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen-32B puts itself at a disadvantaged position. In other words, models’ parameter count affects knowledge more than reasoning. When a compression method aggressively removes the weights of an LRM , it is expected that the model’s knowledge will be more seriously affected. This phenomenon can also be seen on our quantized and pruned models. Since quantization preserves parameter count and our analysis above shows that the quantized models still retain competitive reasoning capability, it is not surprising that even the 1.58-bit model outperforms other distilled and pruned LRMs on MuSiQue. In addition, we notice that pruned R1-Distill-Llama collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024. Since MuSiQue is difficult due to its knowledge requirement, we see that pruning hurts LRMs’ knowledge memorization more than quantization.
Based on the finding that parameter count has a much greater impact on knowledge than reasoning, we recommend quantization as the LRMs compression method when some levels of models’ parametric knowledge are desired. Since both distillation and pruning involve removing weights, we recommend to choose a large student model or a low sparsity when the end task requires knowledge memorization. On the other hand, a model with strong reasoning capability does not need to be extremely large. As Qwen-32B is a stronger reasoner than LLaMA-70B, smaller models also have the potential to outperform larger ones on various reasoning tasks.
5.4 Test-time Compute
We study the behavior of R1 and its compressed variants by measuring their test-time compute. Table 4 shows the analysis of test-time compute when we select the shortest and longest 30% of responses output by each model on each benchmark. We observe that shorter model outputs consistently yield better performance across three reasoning benchmarks, with only an exception on MuSiQue. Regardless of whether an LRM is compressed, if it generates significantly more tokens for a question than other questions in the same dataset, the answer is likely to be incorrect. We exclude Temporal here, because many of the compressed R1 models achieve close to 100% accuracy. The length ratios between the longest and the shortest 30% are typically greater than 4, which indicates nontrivial length differences among model outputs.
Similarly, Table 5 shows the selection of the shortest and the longest 20% of model responses. Compared to Table 4, as we move toward the extreme, the performance gap between the shortest and longest responses becomes even larger. Both R1 and its compressed variants achieve higher scores when they spend less compute during test time. After manually checking some long responses, we notice that longer outputs tend to be more verbose and involve more backtracking in reasoning. This finding demonstrates the need to reduce verbosity to improve reasoning performance, which aligns with recent research (Zhang et al., 2024b).
5.5 Case Study
As discussed in Section 5.2, pruned models collapse on all benchmarks at certain sparsity levels. We identify a common phenomenon when a model collapses: it repeatedly generates a sentence or a chunk until reaching the maximum generation length. We show two examples of this phenomenon in Figure 2. For brevity, we omit the beginning and the end of outputs.
Figure 2: Two examples of the case when a model collapses and keeps repeating itself. These are two outputs (for a FOLIO question) from R1-Distill-Qwen at either 70% or 80% sparsity levels.
In both examples, the pruned models are repeating themselves without pushing their reasoning chains forward, which is a signal of model collapse. When R1-Distill-Qwen is pruned to 70% sparsity, we see that it can still organize a few sentences (e.g., a chunk) to repeat. But when it is pruned to 80% sparsity, it only repeats a simple sentence. This decline of linguistic capability is common when models are pruned to high sparsities. Therefore, aggressively pruning LRMs requires careful consideration.
6 Conclusion and Future Directions
In this paper, we benchmark compressed LRMs on four complex reasoning tasks: AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue. We evaluate models that are quantized, distilled, and pruned. Through MuSiQue, we find that parameter count has a much greater impact on LRMs’ knowledge than reasoning capability, which can potentially guide compression decision. Based on our study of test-time compute, we also find that shorter outputs of R1 and compressed LRMs generally achieve better performance than longer outputs. This finding highlights the need to reduce verbosity in reasoning chains.
Since compression on LRMs is relatively underexplored, future directions involve designing novel methodologies to specifically compress LRMs. The evaluation of compressed LRMs should also place greater emphasis on diverse reasoning tasks. Since DeepSeek-R1 has already been compressed through quantization and distillation, the question remains how to prune such a large model.
Appendix A Additional Details of Reasoning Benchmarks
AIME 2024555https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 (parts I and II) represents top match challenges, and its answers are integers. FOLIO [^6] https://huggingface.co/datasets/yale-nlp/FOLIO requires logical deductions to determine whether the provided conclusion is true, false, or uncertain based on premise. In Temporal Sequences [^7] https://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/temporal_sequences.json, models are asked to use a provided timeline to determine what time a person might be free to perform another activity. Since each of its questions comes with four options, we expect our models to output the index (the letter) of the selected option. The answers of MuSiQue questions are in a few words.