### v2

# When Reasoning Meets Compression: Understanding the Effects of LLMs Compression on Large Reasoning Models

Abstract
Compression methods, including quantization, distillation, and pruning, improve the computational efficiency of large reasoning models (LRMs). However, existing studies either fail to sufficiently compare all three compression methods on LRMs or lack in-depth interpretation analysis. In this paper, we investigate how the reasoning capabilities of LRMs are compromised during compression, through performance benchmarking and mechanistic interpretation. To uncover the effects of compression on reasoning performance, we benchmark quantized, distilled, and pruned DeepSeek-R1 models on four reasoning datasets (AIME 2024, FOLIO, Temporal Sequences, and MuSiQue). To precisely locate compression effects on model weights, we adapt difference of means and attribution patching techniques, focusing on the activation of every linear component in compressed LRMs, to interpret fine-grained causal relationships between weights and various reasoning capabilities. This fine-grained interpretation addresses a fundamental question of compression: which weights are the most important for reasoning? Overall, we find dynamically quantized 2.51-bit R1 reaches close-to-R1 performance. With empirical verification, we present three main findings that generalize across both Llama and Qwen: (1) Weight count has a greater impact on LRMs’ knowledge memorization than reasoning, highlighting the risks of pruning and distillation; (2) The MLP up projection in the final layer of distilled LRMs is one of the most important components, offering a new perspective on locating critical weights — a fundamental problem in model compression; and (3) Current quantization methods overly compress the final-layer modules and MLP gate projections, so protecting just 2% of all weights that are excessively compressed can raise average accuracy by 6.57%, greatly surpassing the state-of-the-art.
1 Introduction
Large reasoning models (LRMs) such as DeepSeek-R1 (Guo et al., 2025) excel at complex reasoning tasks. However, due to their large sizes, deploying them can be costly and even infeasible for individuals, which hinders AI democratization. Compression methods including quantization, distillation, and pruning reduce computational resources (e.g., GPU memory and disk space). Representative quantization techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization AWQ (Lin et al., 2024), and post-training quantization GPTQ (Frantar et al., 2022). Current distillation involves black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings. Representative pruning techniques include unstructured (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
However, existing works do not sufficiently study the performance of compression method on LRMs (Liu et al., 2025; Srivastava et al., 2025; Feng et al., 2025). Although current quantization and pruning methods claim to preserve the performance of general-purpose LLMs, benchmarking both of them on LRMs with more reasoning-intensive datasets helps compare their collapse point. Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider distillation effect on knowledge and reasoning (Huang et al., 2024; Agarwal et al., 2024). Another research gap is the lack of interpretability of compression effects on LRMs. It is necessary to interpret how compression methods affect LRMs, as such analysis can reveal existing bottlenecks and provide guidance for future compression research.
Figure 1: An overview of our pipeline. Left: We benchmark compressed R1 variants on various reasoning tasks. Right: By computing weight importance towards a specific reasoning behavior (a dot product of the steering vector and gradients with respect to an LRM’s activations), we study the compression effects on individual weight. We empirically verify our findings on weight importance by selectively quantizing or protecting a module to test its importance.
Therefore, due to the lack of compression works on LRMs, we study this fundamental research question: How are the reasoning capabilities of LRMs compressed during compression? We answer it from two perspectives — performance benchmarking and mechanistic interpretation. We first benchmark compressed DeepSeek-R1 on various reasoning tasks to investigate how model compression affects performance. We test dynamic quantization (Daniel Han & team, 2023), distillation with supervised fine-tuning (SFT) (Guo et al., 2025), SparseGPT (Frantar & Alistarh, 2023), AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) on R1 (or distilled R1). Then, we apply mechanistic interpretability to quantify weight contribution towards four core reasoning capabilities of LRMs: backtracking, uncertainty estimation, example testing, and adding knowledge. By focusing on the activation of every linear component in compressed LRMs, we adapt difference of means (Arditi et al., 2024) to extract steering vectors and attribution patching (Syed et al., 2023) to compute weight importance. Unlike previous analysis (Venhoff et al., 2025) that only measures layer-wise weight contribution, our weight importance scores offer more fine-grained interpretation of weight contribution, addressing the fundamental compression question of locating important weights. By comparing weight importance scores between compressed LRMs and original LRMs, we quantify the effects of distillation and quantization on model weights [^1] Our interpretation code is at https://github.com/psunlpgroup/Compression-Effects.. Our analysis framework is shown in Figure 1.
With empirical verification, our key findings are summarized below:
  • Weight count has a greater impact on LRMs’ knowledge memorization than their reasoning capabilities, highlighting the compression effects of pruning and distillation. Thus, both distillation and pruning are discouraged when tasks require LRMs’ parametric knowledge.
  • The mlp.up_proj in the final layer of R1 distilled models emerges as one of the most important model components, addressing a core concern in pruning and quantization literature: identifying critical weights. Quantizing only this matrix to 3-bit reduces the average accuracy by 16.3%.
  • Final-layer modules, along with the mlp.gate_proj of R1 distilled Llama and Qwen, are overly compressed by popular quantization methods, highlighting the need for greater attention to preserving their weight precision. A successful protection of only final-layer MLP modules could raise average accuracy by 6.57%, with gains of up to 23.17% over the state-of-the-art quantization.
2 Problem Formulation
2.1 Background
As discussed in Section 1, compression on LRMs (not LLMs) is relatively underexplored. We conduct a thorough literature review in Appendix B.
Bottlenecks on evaluation. Few quantization or pruning methodologies have sufficiently demonstrated effectiveness on LRMs. Current works evaluate quantization and pruning performance primarily using perplexity and simple end tasks, such as the EleutherAI evaluation harness (Gao et al., 2024) and commonsense reasoning. However, quantized or pruned LRMs should be assessed on more complex reasoning tasks with varying difficulty levels. For distillation, although recent works tend to test on more challenging reasoning tasks (compared to other compression literature) such as GSM8K (Cobbe et al., 2021), it is unclear how the compression of LRMs affects models’ parametric knowledge and reasoning capability. Some of them do not comprehensively select diverse reasoning benchmarks (Agarwal et al., 2024). Our benchmarking aims to address these bottlenecks.
Bottlenecks on in-depth analysis. The lack of interpretability of compression effects on LRMs is a key bottleneck of in-depth analysis for compressed LRMs. Being able to interpret the difference between original and compressed LRMs offers a new way to analyze the effects of compression. As a result, better compression approaches can be developed. A recent work (Venhoff et al., 2025) interprets several R1 distilled LRMs, but their focus is not on understanding compression effects.
Recent efforts. Recent benchmarking (Liu et al., 2025) and survey (Feng et al., 2025; Srivastava et al., 2025) papers have begun to evaluate compressed LRMs on more complex reasoning datasets, but they all lack in-depth interpretation of compression effects and do not comprehensively compare different compression strategies. As for compressed LRMs, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights. DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation. Our interpretation analysis aims to demystify the effects of LLMs compression on LRMs, providing a systematic understanding of existing compressed LRMs.
2.2 Mechanistic Interpretation
For our interpretation analysis, we target four core reasoning behaviors following an existing work (Venhoff et al., 2025): backtracking, uncertainty estimation, example testing, and adding knowledge. We prompt GPT-4o to locate token sequences of each behavior from the output tokens of our LRMs. To interpret different compression strategies, we adapt difference of means and attribution patching by computing the activation of every linear module in each layer. This allows us to compute the causal relationship between each weight matrix and our target reasoning behaviors.
Difference of Means. To compute the numerical representation in activation space of each reasoning behavior, we adapt difference of means method (Venhoff et al., 2025; Arditi et al., 2024) to extract the steering vector $\mathbf{u}_{m\ell}^{c}$ for each linear module $m$ at layer $\ell$ for behavior $c$:
$\mathbf{u}_{m\ell}^{c}=\frac{1}{\lvert\mathcal{D}_{+}\rvert}\sum_{s_{i}^{c}\in\mathcal{D}_{+}}\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})-\frac{1}{\lvert\mathcal{D}_{-}\rvert}\sum_{s_{j}\in\mathcal{D}_{-}}\overline{\mathbf{a}}_{m\ell}(s_{j}),\quad\text{with}\quad\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})=\frac{1}{\lvert s_{i}^{c}\rvert}\sum_{t\in s_{i}^{c}}\mathbf{a}_{m\ell}(t)$
where $s_{i}^{c}$ denotes the token sequence corresponding to a specific reasoning behavior $c$ along with its five preceding tokens as output by an LRM, $s_{j}$ is the token sequence of the entire LRM output (prompt and output tokens), $\mathcal{D}_{+}$ is the set of output instances containing at least one token sequence labeled with $c$, $\mathcal{D}_{-}$ is the set of all output instances, $\mathbf{a}_{m\ell}(t)$ is the activation of module $m$ at layer $\ell$ at token $t$, $\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})$ is the average of $\mathbf{a}_{m\ell}(t)$ across all tokens in $s_{i}^{c}$, and similarly, $\overline{\mathbf{a}}_{m\ell}(s_{j})$ is the average of $\mathbf{a}_{m\ell}(t)$ across all tokens in $s_{j}$. We then normalize $\mathbf{u}_{m\ell}^{c}$ to $\tilde{\mathbf{u}}_{m\ell}^{c}$: $\tilde{\mathbf{u}}_{m\ell}^{c}=\mathbf{u}_{m\ell}^{c}\cdot\frac{\left\lVert\overline{\mathbf{a}}_{m\ell}^{\mathrm{all}}\right\rVert_{2}}{\left\lVert\mathbf{u}_{m\ell}^{c}\right\rVert_{2}}$ where $\overline{\mathbf{a}}_{m\ell}^{\mathrm{all}}$ denotes the mean activation across all tokens in $\mathcal{D}_{-}$.
Attribution Patching. To find the causally relevant LRMs components with respect to each reasoning behavior, we adapt attribution patching (Syed et al., 2023) method to compute the importance score $\mathbf{I}_{m\ell}^{c}$ of each linear module.
$\mathbf{I}_{m\ell}^{c}\approx\frac{1}{\lvert\mathcal{D}_{+}\rvert}\left|\sum_{s_{i}^{c}\in\mathcal{D}_{+}}\big(\tilde{\mathbf{u}}_{m\ell}^{\,c}\big)^{\top}\,\frac{\partial}{\partial\mathbf{a}_{m\ell}}\,\mathcal{L}\!\left(s_{i}^{c}\right)\right|$
where $\mathcal{L}\!\left(s_{i}^{c}\right)$ is the cross-entropy loss of $s_{i}^{c}$. A higher $\mathbf{I}_{m\ell}^{c}$ means a stronger causal relationship between $c$ and the linear module $m$ at layer $\ell$, helping us locate the most important weights responsible for reasoning capabilities (a fundamental problem for quantization and pruning works).
2.3 Decoding Compression Effects
To decode compression effects, we compute the relative importance $\mathbf{RI}_{m\ell}^{c}$ of each weight matrix ($\mathbf{I}_{m\ell}^{c}$ divided by $\sum_{m}\sum_{\ell}\mathbf{I}_{m\ell}^{c}$) and track how it changes because of compression (importance shift). Specifically, we measure the change of $\mathbf{RI}_{m\ell}^{c}$ from R1 distilled Llama-8B to original meta-llama/Llama-3.1-8B to understand distillation effect (Section 4). Likewise, the importance shift from the R1 distilled models to their quantized versions indicates quantization effect (Section 5). For the R1 distilled models, we also compute the $\mathbf{I}_{m\ell}^{c}$ of their weights to complement our findings on the distillation effect.
We hypothesize that the importance shift should be minimal in the ideal case, as a compressed LRM should remain as close as possible to its original counterpart (the more reasoning-capable model). When visualizing the importance shift from an LRM to its compressed variant (or from a distilled model to its backbone), we only consider decreases in $\mathbf{RI}_{m\ell}^{c}$. By definition, the relative importance of each weight matrix is normalized to sum to one, so any increase in relative importance necessarily compensates for decreases elsewhere. Since it is more informative to track cases where the $\mathbf{RI}_{m\ell}^{c}$ of a more reasoning-capable model decreases (e.g., when the reasoning capability of a weight matrix is diminished), we set all increases in relative importance to zero.
2.4 Scope
We study three major LLMs compression paradigms, distillation, quantization, and pruning, making our scope comprehensive enough for investigating the effects of diverse compression methods. For distillation, we select four R1 distilled models: DeepSeek-R1-Distill Llama-70B, Qwen-32B, Llama-8B, and Qwen-7B. For quantization, we select 2.51-, 1.73-, and 1.58-bit models by Unsloth [^2] https://huggingface.co/unsloth/DeepSeek-R1-GGUF as the choices of quantized R1 due to their popularity. We also evaluate AWQ Lin et al. (2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) as reproducible state-of-the-art quantization methods designed for relatively smaller LLMs (e.g., the R1 distilled models). Specifically, we use all four methods to perform 4-bit quantization, and use GPTQ, GPTAQ, and ANY3 for 3-bit quantization as well, since many AWQ implementations do not support 3-bit. For pruning, we run SparseGPT Frantar & Alistarh (2023) on our two largest distilled models. We run interpretaion analysis on linear modules of all layers within LRMs.
2.5 Evaluation Setup
We select four reasoning datasets with varying levels of difficulty: AIME 2024 (Mathematical Association of America, ) for mathematical reasoning, FOLIO (Han et al., 2024) for logical reasoning, Temporal Sequences of BIG-Bench Hard (Suzgun et al., 2022) for temporal reasoning, and MuSiQue (Trivedi et al., 2022) for multihop reasoning. Since MuSiQue requires knowledge memorization besides multihop reasoning, we follow a closed-book setting (directly prompting LRMs to get final answers) to evaluate both reasoning and knowledge retention capabilities. Additional details of benchmarks, along with Table 5 that shows their statistics, are specified in Appendix C.
Accuracy metric is used for AIME 2024, FOLIO, and Temporal Sequences. We adopt exact match (EM) and F1 for MuSiQue. For each model (except R1 and those dynamically quantized LRMs), we run it three times and report its average scores to mitigate performance variability. Implementation details are in Appendix D.
3 Compression Effects on Reasoning Performance
3.1 Overall Performance
The overall performance of R1 and its compressed variants are in Table 1. We show the performance of pruned R1-Distill-Llama-70B and R1-Distill-Qwen-32B under 50% sparsity in Table 1, as it is the default sparsity level of current works (Zhang et al., 2024a; Sun et al., 2023). Additional analysis of test-time compute and model collapse behavior are in Appendices G and H.
Table 1: Benchmark performance of R1 and its compressed variants. All four benchmark scores are averaged over three passes, except the rows marked with . Avg denotes the average scores shown in AIME 2024, FOLIO, and Temporal columns. We segment this table based on model families and mark the highest scores within each model family in bold.
Models Accuracy
Model #Param Compression AIME 2024 FOLIO Temporal Avg MuSiQue (EM, F1)
DeepSeek-R1 671B - 73.3 76.4 99.6 83.1 (17.0, 27.51)
DeepSeek-R1 671B 2.51-bit 76.7 77.8 100.0 84.8 (17.0, 24.43)
DeepSeek-R1 671B 1.73-bit 66.7 78.3 99.6 81.5 (15.0, 22.11)
DeepSeek-R1 671B 1.58-bit 66.7 75.4 94.0 78.7 (14.0, 22.34)
R1-Distill-Llama 70B Distillation 65.6 79.8 99.9 81.8 (13.3, 21.57)
R1-Distill-Llama 70B Distillation & 50% sparse 23.3 71.6 97.6 64.2 (6.7, 13.49)
R1-Distill-Llama 70B Distillation & 4-bit AWQ 63.4 78.5 99.3 80.4 (10.7, 19.23)
R1-Distill-Llama 70B Distillation & 4-bit GPTQ 66.7 77.0 99.9 81.2 (10.3, 18.17)
R1-Distill-Llama 70B Distillation & 4-bit GPTAQ 64.4 78.8 99.6 80.9 (12.0, 21.57)
R1-Distill-Llama 70B Distillation & 3-bit GPTQ 46.7 71.8 99.3 72.6 (4.7, 11.92)
R1-Distill-Llama 70B Distillation & 3-bit GPTAQ 54.4 77.3 99.7 77.1 (5.7, 13.21)
R1-Distill-Qwen 32B Distillation 64.4 82.3 99.9 82.2 (2.7, 10.95)
R1-Distill-Qwen 32B Distillation & 50% sparse 25.6 75.1 97.9 66.2 (2.3, 9.01)
R1-Distill-Qwen 32B Distillation & 4-bit AWQ 67.8 82.3 99.1 83.1 (3.3, 10.28)
R1-Distill-Qwen 32B Distillation & 4-bit GPTQ 68.9 80.6 99.6 83.0 (4.0, 11.78)
R1-Distill-Qwen 32B Distillation & 4-bit GPTAQ 63.3 81.5 99.7 81.5 (2.7, 11.88)
R1-Distill-Qwen 32B Distillation & 4-bit ANY4 68.9 78.0 99.7 82.2 (5.7, 12.68)
R1-Distill-Qwen 32B Distillation & 3-bit GPTQ 44.4 74.2 98.9 72.5 (4.0, 11.55)
R1-Distill-Qwen 32B Distillation & 3-bit GPTAQ 45.6 77.5 99.5 74.2 (2.3, 9.18)
R1-Distill-Qwen 32B Distillation & 3-bit ANY3 53.3 82.6 99.9 78.6 (3.7, 10.27)
R1-Distill-Llama 8B Distillation 42.2 71.9 81.5 65.2 (0.0, 4.43)
R1-Distill-Llama 8B Distillation & 4-bit AWQ 47.8 68.0 84.0 66.6 (0.3, 5.05)
R1-Distill-Llama 8B Distillation & 4-bit GPTQ 42.2 66.2 65.9 58.1 (0.3, 4.68)
R1-Distill-Llama 8B Distillation & 4-bit GPTAQ 40.0 66.4 69.3 58.6 (0.0, 3.73)
R1-Distill-Llama 8B Distillation & 4-bit ANY4 41.1 68.5 88.7 66.1 (0.0, 3.54)
R1-Distill-Llama 8B Distillation & 3-bit GPTQ 11.1 65.0 67.3 47.8 (0.0, 2.89)
R1-Distill-Llama 8B Distillation & 3-bit GPTAQ 7.8 65.5 57.2 43.5 (0.0, 3.45)
R1-Distill-Llama 8B Distillation & 3-bit ANY3 3.3 50.1 34.9 29.4 (0.7, 2.35)
R1-Distill-Qwen 7B Distillation 46.7 78.0 75.6 66.8 (0.0, 3.57)
R1-Distill-Qwen 7B Distillation & 4-bit AWQ 46.6 75.5 74.9 65.7 (0.0, 3.14)
R1-Distill-Qwen 7B Distillation & 4-bit GPTQ 38.9 72.9 70.3 60.7 (1.0, 4.27)
R1-Distill-Qwen 7B Distillation & 4-bit GPTAQ 47.8 74.4 67.7 63.3 (0.0, 3.96)
R1-Distill-Qwen 7B Distillation & 4-bit ANY4 47.8 75.6 77.1 66.8 (0.0, 3.05)
R1-Distill-Qwen 7B Distillation & 3-bit GPTQ 17.8 65.7 31.7 38.4 (0.0, 3.12)
R1-Distill-Qwen 7B Distillation & 3-bit GPTAQ 24.4 64.5 48.7 45.9 (0.0, 3.06)
R1-Distill-Qwen 7B Distillation & 3-bit ANY3 32.2 69.3 30.1 43.9 (0.0, 3.89)
Comparing Compression Strategies. In Table 1, the 2.51-bit R1 achieves the highest average accuracy overall, since it has the smallest compression ratio. Both R1 distilled Llama-70B and Qwen-32B reach close-to-R1 accuracy scores. On MuSiQue, the 2.51-bit R1 also achieves performance close to original R1. Therefore, 2.51-bit R1 has the best overall performance than other compression strategies. Although R1 may be over-parameterized, a compression method with a smaller ratio can still offer advantages over methods with higher compression ratios. In contrast, pruning only 50% of the weights causes significant degradation, rendering the pruned LRMs unusable. Thus, we do not interpret the effect of pruning in the later sections. As for all distillation-only models, Qwen delivers stronger reasoning performance than Llama (Appendix E).
Comparing Benchmark Difficulties. Comparing the scores using R1 distilled Llama-70B on AIME 2024, FOLIO, and Temporal, we see the largest score decrease on AIME 2024. This indicates that AIME 2024 is more challenging than the other two accuracy-based benchmarks. MuSiQue is also difficult in terms of knowledge requirement, because its scores in Table 1 are much lower than RAG (retrieval-augmented generation) setup (Zhang et al., 2025). This suggests that existing LRMs lack sufficient knowledge for knowledge-intensive tasks, making RAG a more suitable approach.
Takeaway 3.1 for Overall Performance Considering over-parameterization, methods with smaller compression ratios can still offer advantages over those with higher compression ratios. Regardless of whether compression is applied, LRMs lack sufficient knowledge for knowledge-intensive tasks.
3.2 Collapse Point
We investigate whether LRMs degrade as they undergo increasing levels of compression. In Table 1, the performance of dynamically quantized LRMs steadily declines as we move from 2.51 to 1.58-bit, but we do not observe a clear collapse point. All 4-bit AWQ, GPTQ, GPTAQ, and ANY4 reach performance similar to their unquantized counterparts, which shows the effectiveness of existing 4-bit quantization on LRMs. However, 3-bit GPTQ, GPTAQ, and ANY3 display signs of collapse, indicating bottlenecks of current 3-bit quantization. GPTAQ and ANY4 are newer than AWQ and GPTQ, and they achieve similar performance on 4-bit LRMs. Based on average accuracy, it is noteworthy that GPTAQ surpasses GPTQ on 3-bit LRMs for three out of four distilled models. Regarding distillation, R1 distilled Llama-8B and Qwen-7B achieve the lowest accuracy among all distillation-only models. Only R1 distilled Llama-70B yields decent MuSiQue scores.
Table 2: Performance of two distilled models under various sparsity levels. We report the one-pass scores for all models in this table.
Models Accuracy
Model #Param Sparsity AIME 2024 FOLIO Temporal Avg MuSiQue (EM, F1)
R1-Distill-Llama 70B 0% 63.3 78.8 100.0 80.7 (13.0, 21.80)
R1-Distill-Llama 70B 10% 60.0 81.3 99.6 80.3 (12.0, 21.69)
R1-Distill-Llama 70B 30% 63.3 79.3 99.6 80.7 (14.0, 21.40)
R1-Distill-Llama 70B 40% 56.7 73.9 98.8 76.8 (6.0, 13.79)
R1-Distill-Llama 70B 50% 26.7 70.9 97.2 64.9 (6.0, 12.75)
R1-Distill-Llama 70B 60% 0.0 65.0 95.6 53.5 (0.0, 6.42)
R1-Distill-Llama 70B 70% 0.0 49.8 15.6 21.8 (0.0, 2.23)
R1-Distill-Llama 70B 80% 0.0 11.8 12.4 8.1 (0.0, 0.94)
R1-Distill-Qwen 32B 0% 66.7 82.3 100.0 83.0 (1.0, 9.38)
R1-Distill-Qwen 32B 10% 70.0 81.3 100.0 83.8 (5.0, 13.19)
R1-Distill-Qwen 32B 30% 56.7 81.3 100.0 79.3 (1.0, 10.47)
R1-Distill-Qwen 32B 40% 53.3 78.3 100.0 77.2 (2.0, 10.16)
R1-Distill-Qwen 32B 50% 30.0 75.4 96.0 67.1 (3.0, 9.29)
R1-Distill-Qwen 32B 60% 0.0 65.0 87.2 50.7 (0.0, 4.13)
R1-Distill-Qwen 32B 70% 0.0 32.5 19.6 17.4 (0.0, 1.72)
R1-Distill-Qwen 32B 80% 0.0 8.7 2.0 3.6 (0.0, 1.29)
Table 2 displays performance of our two distilled models under various sparsity levels. Comparing distilled models with their sparsified variants, we find the precise collapse points of our pruned LRMs. Interestingly, their collapse points correlate to the benchmark difficulty. For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half. However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024. The correlation between collapse point and benchmark difficulty can also be seen on the sparsified Qwen.
Takeaway 3.2 for Collapse Point Collapse point correlates with benchmark difficulty. On hard benchmarks, 3-bit quantization and pruning with 50% sparsity or higher still have substantial room for improvement.
3.3 Compression Impact on Knowledge and Reasoning
In Table 2, although Qwen demonstrates stronger reasoning capabilities than Llama, it has significantly lower EM and F1 scores on MuSiQue. Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen puts itself at a disadvantaged position. In other words, models’ parameter count affects knowledge more than reasoning. In addition, we notice pruned R1-Distill-Llama-70B collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024. This shows pruning hurts LRMs’ knowledge memorization more than quantization. When a compression method aggressively removes the weights of an LRM, it is expected that the model’s knowledge will be more seriously affected. This phenomenon can also be seen on our dynamically quantized models in Table 1. Since quantization preserves parameter count and our analysis above shows that many quantized models still retain competitive reasoning capability, quantization is recommended on knowledge-intensive tasks.
Takeaway 3.3 for Compression Impact on Knowledge and Reasoning Pruning and distillation compress knowledge retention more than reasoning capabilities.
4 Distillation Effect on Weights
To study the effect of distillation on weights, we compute $\mathbf{I}_{m\ell}^{c}$ of two distilled R1 models and further measure the change of $\mathbf{RI}_{m\ell}^{c}$ as discussed in Section 2.3.
4.1 Locating Important Weights
The left part of Figure 2 presents the weight importance of R1 distilled Llama-8B in four heatmaps, each corresponding to a reasoning behavior. We observe that the final layer houses several most important linear modules across all four behaviors, with the highest value located at up_proj. Therefore, the up_proj in the final layer (32_up) stands out as the most important component.
Interestingly, this finding generalizes to R1 distilled Qwen-7B, as we also observe this up_proj outlier in the final layer of Qwen in Figure 4. Notably, our finding complements a recent analysis (Shao & Wu, 2025), which claims the most important module for reasoning is o_proj. Since identifying important weights is a core research problem of compression methodologies, our finding is valuable for future works.
Takeaway 4.1 for Locating Important Weights Distillation makes up_proj in the final layer as the most important module for reasoning behaviors, as observed in both R1 distilled Llama and Qwen models.
Figure 2: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Llama-8B (left) and change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Llama-3.1-8B (right). Each heatmap displays scores of importance (or importance shift) of every module at each layer, providing a fine-grained analysis of weight contributions to the corresponding reasoning capability. On the right, increases in $\mathbf{RI}_{m\ell}^{c}$ are set to 0, as they only offset decreases elsewhere as discussed in 2.3. Every cluster of 4 side-by-side heatmaps (including those displayed below) follow the same scaling to show the precise magnitude of each weight module.
Figure 3: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant.
Table 3: Accuracy after selectively quantizing a single component of R1 distilled Llama-8B (e.g., 1_up means to only quantize the up_proj in the first layer) to 3-bit. Ranking of a component is based its $\sum_{c}\mathbf{I}_{m\ell}$, so “2nd col” refers to second place within its column across all four heatmaps (each column consists of 7 linear modules of a layer). “1st overall” means the global highest ranking.
Quantized Component Rank AIME 2024 FOLIO Temporal Avg
32_up 1st overall 20.0 63.1 63.6 48.9
32_gate 2nd col 33.3 62.1 67.2 54.2
32_v last col 43.3 68.0 79.6 63.6
31_up 2nd row 33.3 70.0 64.4 55.9
1_up last row 6.7 64.5 80.4 50.5
4.2 Validating Importance Scores
We validate Section 4.1 by applying 3-bit round-to-nearest quantization to either 32_up or a component sharing its column or row in the heatmaps, then measuring the resulting accuracy drop (Table 3). The more important a component is, the greater the accuracy drop when it is quantized. Specifically, we select four additional component candidates: the second- and last-ranked modules among the seven in the final layer, and the second- and last-ranked layers across all 32 layers of the up-projection. We see that 32_up yields the lowest average accuracy, which clearly demonstrates the validity of our findings in Section 4.1. It is quite salient that quantizing only this matrix (merely 0.7% of all weights) reduces the average accuracy by 16.3%. The component rank generally correlates with the accuracy drop, except for 1_up, which incurs the lowest accuracy on AIME 2024.
4.3 Importance Shift via Distillation
Since R1-Distill-Llama-8B is fine-tuned based on Llama-3.1-8B, we compute the change of $\mathbf{RI}_{m\ell}^{c}$ to visualize distillation effect in the right part of Figure 2 (for Llama) and 5 (for Qwen). Both parts of Figure 2 exhibit similar patterns (e.g., most outliers are in the final layer), indicating that the important weights of the distilled model are primarily the result of distillation with SFT, while the original Llama’s weight values play little role in shaping its reasoning capabilities. Thus, distillation effect is quite powerful in transforming a non-reasoning LLM into an LRM. For Qwen, Figures 4 and 5 also show similar patterns, so the utility of distillation generalizes to Qwen as well.
Takeaway 4.3 for Importance Shift via Distillation Important weights of the R1 distilled models are mainly the result of the distillation effect.
5 Quantization Effect on Weights
For quantization effect on weights, we analyze the decrease of importance shift during quantization.
5.1 Locating Quantization Effect
We show heatmaps to visualize the importance shift from R1 distilled Llama-8B to its 4-bit AWQ quantized variant in Figure 3. Across all four heatmaps, we observe a reduction in the significance of the gate projections in the middle layers (e.g., layer 9 to 23), suggesting that AWQ may overly compress these modules. Moreover, most linear modules in the final layer are compressed to the greatest extent, which shows the drawback of AWQ. Since 32_up is the most important module as discussed in Section 4.2 and its importance shift is little for uncertainty estimation and example testing capabilities, AWQ successfully preserves its significance on these two behaviors. However, for backtracking and adding knowledge capabilities, AWQ is not effective at maintaining its importance.
In Figure 6, we visualize the importance shift from R1 distilled Qwen-7B to its 4-bit AWQ quantized version. We also see a shift in the importance of the gate projections on Qwen, but this shift mainly occurs in the early layers (e.g., layer 1 to 10). On Qwen, AWQ does not preserve the importance of 32_up across all four reasoning capabilities, and it also overly compresses 32_k on two capabilities.
As another popular method, we interpret the effect of 4-bit GPTQ in Figure 7. On R1 distilled Llama-8B, we observe similar quantization effect as AWQ, since GPTQ also overly compresses final-layer modules and the gate projections in the middle layers. Their commonality demonstrates the generality of the bottlenecks we identified in existing quantization methods.
Takeaway 5 for Quantization Effect on Weights State-of-the-art quantization methods fail to preserve the importance of the MLP gate projections and the final layer, which is a key bottleneck of performance improvement.
5.2 Validating Quantization Effect
Table 4: Performance of 3-bit AWQ and selectively protecting the MLP modules in the final layer.
Model Compression Full-Precision Anywhere? AIME 2024 FOLIO Temporal Avg MuSiQue
R1-Distill-Llama-8B 3-bit AWQ - 10.0 59.6 68.4 46.0 (0.0, 3.50)
R1-Distill-Llama-8B 3-bit AWQ Final-layer MLP 16.7 67.0 74.0 52.57 (1.0, 3.62)
To validate our findings about the bottleneck of current quantization, we design a simple protection mechanism using a mixed precision fashion. We run two versions of 3-bit AWQ in Table 4. In the first version, we run AWQ with their default calibration data. Since we know AWQ overly compresses the MLP modules in the final layer, we then choose to protect them by changing their quantized weights to their original values in 16-bit. If they are truly important yet not well protected by AWQ, our protection mechanism should offer a significant improvement. Based on the discussion in Sections 3.1 and 3.2, we perform 3-bit quantization, since it will be the focus of future works.
We see our selective protection boosts 3-bit AWQ on all benchmarks, with an average accuracy improvement of 6.57%. This is particularly significant given that only about 2% of all weights remain in 16-bit. This mixed precision model outperforms all 3-bit quantization baselines in Table 1 by at least 4.77% in average accuracy, with gains of up to 23.17%. Therefore, our findings are demonstrated with an indication of substantial room for further improvement. Note that our protection provides relatively marginal increase on MuSiQue, as the weight count stays the same (Section 3.3).
6 Conclusion and Future Directions
We study the effects of LLMs compression on LRMs and present key findings for further imrpving LRMs compression. Future compression works are encouraged to consider the protection of MLP up projection in the final layer. The excessive compression of current quantization methods on MLP gate projections and final-layer modules highlights the need for preserving these weight modules.
Appendix A Use of LLMs
To improve the overall clarity of our writing, we used ChatGPT-4o and ChatGPT-5 via OpenAI’s web interface to polish a small fraction of sentences. LLMs were not used in any steps of the research ideation process. To ensure correctness and precision, we carefully reviewed and adapted all LLM-generated content before incorporating it into our writing.
Appendix B Related Work
Our literature review is conducted over existing compression methodologies (quantization, distillation, and pruning) and recent LRMs.
B.1 Quantization
Quantization reduces the number of bits used to represent LLM weights, thereby lowering their precision (Srivastava et al., 2025). Recent survey (Zhu et al., 2024) categorizes quantization methodologies into quantization-aware training (QAT) and post-training quantization (PTQ). QAT requires retraining of model weights to recover performance loss during quantization while PTQ does not require retraining. Recent QAT includes LLM-QAT (Liu et al., 2024a) that adopts distillation to train a quantized LLM, BitDistiller (Du et al., 2024) that develops a self-distillation approach for the full-precision model to act as the teacher of its low-bit counterpart, BitNet (Wang et al., 2023) that proposes a 1-bit Transformer architecture for training LLMs from scratch, and OneBit (Xu et al., 2024) that quantizes LLM weight matrices to 1-bit from a knowledge transfer perspective.
PTQ is more popular in terms of the number of recent publications, because there is no retraining involved. For example, GPTQ (Frantar et al., 2022) and GPTAQ (Li et al., 2025) are one-shot weight quantization methods that use approximate second-order information, while AWQ (Lin et al., 2024) leverages activation distribution for finding the salient weight channels to skip. Other PTQ methods include weight-activation quantization (Shao et al., 2024; Yao et al., 2022; Liu et al., 2023) and KV cache quantization (Hooper et al., 2024; Liu et al., 2024b).
B.2 Distillation
Distillation involves two settings: black-box and white-box settings. For black-box setting, the teacher model is typically a closed-source LLM and only the outputs of teacher are available for the student model. For white-box setting, both weights and output distribution of the teacher model are available. Existing black-box distillation (Huang et al., 2024; Li et al., 2024b; Ho et al., 2023; Huang et al., 2022; Li et al., 2024a) prompts the teacher model to generate a training dataset for the student to learn. Specifically, researchers have started to distill OpenAI’s O1 model (Huang et al., 2024), which marks the beginning of LRMs compression. White-box distillation allows the student model to learn from teacher’s knowledge representation. Works has been done to align the output distribution (Agarwal et al., 2024; Gu et al., 2024) or the hidden representation (Liang et al., 2023) between teacher and student models.
B.3 Pruning
There are unstructured and structured pruning. For unstructured pruning, individual weights are targeted, which leads to irregular sparsity structure. In contrast, structured pruning involves removing entire network components such as channels or layers (Zhang et al., 2024a). Unstructured pruning usually has better compression performance than structured pruning, while it is easier to achieve inference speedup via structured methods (Zhu et al., 2024). Recent unstructured pruning includes one-shot pruning (Frantar & Alistarh, 2023; Sun et al., 2023), global pruning that makes pruning decisions based on all layers (Bai et al., 2024), and domain-specific pruning (Zhang et al., 2024a). Structured pruning includes gradient-based  (Xia et al., 2024; Ma et al., 2023) and non-gradient-based (Ashkboos et al., 2024) methods.
B.4 LRMs
Trained with reinforcement learning, LRMs extends LLMs with advanced reasoning mechanisms (Besta et al., 2025). Popular closed-source LRMs are OpenAI’s o1-mini, o1 (OpenAI et al., 2024), and o3-mini. Open-source LRMs include DeepSeek-R1 and QwQ-32B-Preview (Team, 2024). Since quantization, white-box distillation, and pruning methods require access to model weights, they are not suitable for closed-source LRMs. Only black-box distillation will work on closed-source models.
Table 5: Dataset statistics of selected reasoning benchmarks.
Size Answer Type Metric Knowledge
AIME 2024 30 Integer Accuracy False
FOLIO 203 True/False/Uncertain Accuracy False
Temporal 250 (A)/(B)/(C)/(D) Accuracy False
MuSiQue 100 A few words (EM, F1) True
Appendix C Additional Details of Reasoning Benchmarks
Table 5 shows the statistics of our selected benchmarks. AIME 2024333https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 (parts I and II) represents top match challenges, and its answers are integers. FOLIO [^4] https://huggingface.co/datasets/yale-nlp/FOLIO requires logical deductions to determine whether the provided conclusion is true, false, or uncertain based on premise. In Temporal Sequences [^5] https://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/temporal_sequences.json, models are asked to use a provided timeline to determine what time a person might be free to perform another activity. Since each of its questions comes with four options, we expect our models to output the index (the letter) of the selected option. Since MuSiQue involves question answering and its answers are in a few words, we adopt exact match (EM) and F1. We randomly sample 100 questions out of 1000 from MuSiQue for our benchmarking analysis.
Appendix D Implementation Details
We run the dynamically quantized models on llama.cpp [^6] https://github.com/ggml-org/llama.cpp based on their requirement. We run all other distilled, pruned, and quantized models on vLLM (Kwon et al., 2023) for its fast inference. In order to comprehensively analyze performance change after compression, we also evaluate R1 on our reasoning benchmarks by using DeepSeek API. Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models during inference: maximum generation length is set to $32768$, temperature is set to $0.6$, and top-p value is set to $0.95$.
We use AutoAWQ [^7] https://github.com/casper-hansen/AutoAWQ as the AWQ implementation for inference due to its speed advantage (vLLM support), while the original AWQ code [^8] https://github.com/mit-han-lab/llm-awq is used to generate the pseudo-quantized R1 distilled Llama-8B for our analysis in Section 5.
We focus on analyzing the effect of compression methods on performance and thus do not consider inference speedup. The reason is that these methods run on different inference platforms, so it is hard to control the consistency of inference optimization across various platforms.
Appendix E Comparing Distilled Models
On accuracy-based benchmarks of Table 1, we see that R1 distilled Qwen-32B delivers an average 0.4% improvement over Llama-70 and R1 distilled Qwen-7B delivers an average 1.6% improvement over Llama-8B. Although these two Qwen models have less weights, Qwen delivers stronger reasoning performance than Llama. This phenomenon aligns with DeepSeek report (Guo et al., 2025). However, R1-Distill-Qwen-32B scores significantly lower than R1-Distill-Llama-70B on MuSiQue, highlighting its worse ability of memorization.
Appendix F Additional Visualization of Weight Importance and Importance Shift
All additional figures in Appendix are thoroughly discussed in the main content. Figure 4 shows the weight importance of DeepSeek-R1-Distill-Qwen-7B across four heatmaps, each corresponding to a specific target reasoning behavior. Figure 5 displays the change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Qwen2.5-Math-7B. To decode the quantization effect on Qwen, Figure 6 shows the change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Qwen-7B to its 4-bit AWQ variant. Similarly, Figure 7 shows the change of $\mathbf{RI}_{m\ell}^{c}$ from R1 distilled Llama-8B to its 4-bit GPTQ quantized variant.
Figure 4: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Qwen-7B.
Figure 5: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Qwen2.5-Math-7B (the backbone model).
Figure 6: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Qwen-7B to its 4-bit AWQ variant.
Figure 7: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit GPTQ quantized variant.
Appendix G Test-time Compute
We study the behavior of R1 and some of its compressed variants by measuring their test-time compute. Table 6 shows the analysis of test-time compute when we select the shortest and longest 30% of responses output by each model on each benchmark. We observe that shorter model outputs consistently yield better performance across three reasoning benchmarks, with only an exception on MuSiQue. Regardless of whether an LRM is compressed, if it generates significantly more tokens for a question than other questions in the same dataset, the answer is likely to be incorrect. We exclude Temporal here, because many of the compressed R1 models achieve close to 100% accuracy. The length ratios between the longest and the shortest 30% are typically greater than 4, which indicates nontrivial length differences among model outputs.
Similarly, Table 7 shows the selection of the shortest and the longest 20% of model responses. Compared to Table 6, as we move toward the extreme, the performance gap between the shortest and longest responses becomes even larger. Both R1 and its compressed variants achieve higher scores when they spend less compute during test time. After manually checking long responses, we notice that longer outputs tend to be more verbose and involve more backtracking in reasoning. This finding demonstrates the need to reduce verbosity to improve reasoning performance, which aligns with recent research (Zhang et al., 2024b).
Table 6: Analysis of test-time compute when selecting the shortest and longest 30% of responses output by each model on each benchmark. “Short” column contains performance scores of the shortest 30% of outputs from a model, while “long” column contains scores of the longest 30% of outputs. We compare the scores between “Short” and “Long” for every model and benchmark, and mark the best scores in bold. “Ratio” column represents the ratio of the average length (in number of tokens) of the longest 30% to that of the shortest 30%.
Models AIME 2024 FOLIO MuSiQue
Model Compression Short Long Ratio Short Long Ratio Short Long Ratio
DeepSeek-R1 - 88.9 33.3 5.3 83.3 63.3 8.0 (30.0, 42.8) (3.3, 10.0) 6.4
DeepSeek-R1 2.51-bit 100.0 33.3 4.9 85.0 71.7 6.5 (33.3, 41.9) (0.0, 6.9) 7.7
DeepSeek-R1 1.73-bit 88.9 22.2 4.4 86.7 73.3 4.9 (30.0, 41.6) (10.0, 23.5) 6.9
DeepSeek-R1 1.58-bit 77.8 44.4 3.8 80.0 65.0 5.0 (30.0, 41.5) (10.0, 16.6) 5.9
R1-Distill-Llama Distillation 88.9 11.1 5.4 80.0 80.0 4.5 (26.7, 38.0) (6.7, 18.4) 4.0
R1-Distill-Llama Distillation & 10% sparse 100.0 0.0 6.6 85.0 78.3 4.9 (20.0, 29.6) (6.7, 13.7) 7.4
R1-Distill-Llama Distillation & 30% sparse 88.9 11.1 5.3 85.0 76.7 5.0 (26.7, 36.9) (3.3, 12.5) 8.8
R1-Distill-Qwen Distillation 100.0 11.1 7.1 86.7 75.0 6.9 (16.7, 24.1) (16.7, 24.7) 7.1
R1-Distill-Qwen Distillation & 10% sparse 88.9 33.3 4.8 83.3 75.0 5.5 (23.3, 36.7) (3.3, 12.8) 8.1
R1-Distill-Qwen Distillation & 30% sparse 88.9 11.1 6.8 86.7 78.3 7.8 (26.7, 40.7) (3.3, 9.1) 8.8
Table 7: Analysis of test-time compute when selecting the shortest and longest 20% of responses output by each model on each benchmark. Refer to Table 6 for the meaning of each column, except that we select shortest and longest 20% instead.
Models AIME 2024 FOLIO MuSiQue
Model Compression Short Long Ratio Short Long Ratio Short Long Ratio
DeepSeek-R1 - 100.0 16.7 7.1 82.5 57.5 10.9 (30.0, 42.4) (5.0, 12.5) 8.5
DeepSeek-R1 2.51-bit 100.0 33.3 6.6 87.5 62.5 8.6 (40.0, 50.3) (0.0, 4.5) 10.7
DeepSeek-R1 1.73-bit 100.0 33.3 5.9 87.5 67.5 6.3 (30.0, 43.7) (5.0, 15.7) 9.3
DeepSeek-R1 1.58-bit 100.0 33.3 5.6 80.0 72.5 6.2 (35.0, 46.3) (15.0, 20.0) 7.6
R1-Distill-Llama Distillation 83.3 16.7 7.0 82.5 75.0 5.7 (35.0, 45.2) (5.0, 9.2) 5.3
R1-Distill-Llama Distillation & 10% sparse 100.0 0.0 9.2 90.0 75.0 6.3 (20.0, 28.5) (5.0, 11.7) 10.7
R1-Distill-Llama Distillation & 30% sparse 100.0 16.7 7.1 90.0 75.0 6.4 (20.0, 31.0) (0.0, 7.3) 13.0
R1-Distill-Qwen Distillation 100.0 0.0 9.8 87.5 75.0 9.0 (20.0, 28.7) (20.0, 24.2) 10.4
R1-Distill-Qwen Distillation & 10% sparse 83.3 33.3 6.1 85.0 75.0 7.2 (30.0, 36.2) (0.0, 8.4) 12.0
R1-Distill-Qwen Distillation & 30% sparse 100.0 0.0 8.9 87.5 75.0 11.0 (30.0, 36.6) (5.0, 11.2) 13.4
Appendix H Case Study
As discussed in Section 3.2, pruned models collapse on all benchmarks at certain sparsity levels. We identify a common phenomenon when a model collapses: it repeatedly generates a sentence or a chunk until reaching the maximum generation length. We show two examples of this phenomenon in Figure 8. For brevity, we omit the beginning and the end of outputs.
Figure 8: Two examples of the case when a model collapses and keeps repeating itself. These are two outputs (for a FOLIO question) from DeepSeek-R1-Distill-Qwen-32B at either 70% or 80% sparsity levels.
In both examples, the pruned models are repeating themselves without pushing their reasoning chains forward, which is a signal of model collapse. When R1-Distill-Qwen-32B is pruned to 70% sparsity, we see that it can still organize a few sentences (e.g., a chunk) to repeat. But when it is pruned to 80% sparsity, it only repeats a simple sentence. This decline of linguistic capability is common when models are pruned to high sparsities. Therefore, aggressively pruning LRMs requires careful consideration.